A two digit number is two less than six timers the sum of its digits. The tens digit is one more than the units digit. Find the number.
Answer by Paul(988)

The tens digit is x
the unit digit is y
x=1+y

10x+y=6(x+y)-2
10x+y=6x+6y-2
Subsitute for y
10(1+y)+y=6(1+y)+6y-2
10+10y+y=6+6y+6y-2
11y+10=4+12y
-y=-6
y=6

x=1+6
x=7

Hence, the 10 digits is 7 and the unit digit is 6.
Paul.